# Two results on the Convex Algebraic Geometry of sets with continuous symmetries

### v1

@@ 3. Equivariant descriptions of invariant spectrahedra
As illustrated by 3.4, one cannot expect that $S_{\mathsf{G}}$ is a spectrahedron, even if $S$ is. The following partial converse to 3.5 shows that if $S_{\mathsf{G}}$ is a spectrahedron, which, moreover, is given by linear matrix inequalities coming from 3.2, then, in fact, $S$ was already $\mathsf{G}$-invariant (assuming $\mathsf{G}$ is connected, cf. 3.6).
@@
By 3.3, we have that $S_{\mathsf{G}}=\{x\in{\mathds{R}}^{n}:\widetilde{M}(x)\succeq 0\}$, so it suffices to show that $S_{\mathsf{G}}=S$, see (3.2). By 2.14 and 2.13, it suffices to prove $\partial S_{\mathsf{G}}\subset\partial S$. Note that $S_{\mathsf{G}}$ is a spectrahedron, since we are assuming $S_{\mathsf{G}}=\{x\in{\mathds{R}}^{n}:\overline{M}(x)\succeq 0\}$. Thus, given $x\in\partial S_{\mathsf{G}}$, the kernel of $\overline{M}(x)$ is nonzero, so there exists $\varphi\neq 0$ in $V$ such that $\widetilde{M}(x)(\varphi,\varphi)=0$, i.e.,
@@
The statement of 3.10 becomes false if we drop the assumption that $\mathsf{G}$ is connected. Indeed, if $\mathsf{G}$ is finite, then ${\mathcal{V}}$ itself is finite-dimensional.
@@ 4. Preliminaries from polar representations
A convex subset $S$ of $\mathfrak{a}\cong{\mathds{R}}^{n}$ that is invariant under $W=S_{n}$ is commonly called symmetric. The corresponding $\mathsf{O}(n)$-invariant convex subset of $V$ is given by the set of symmetric matrices $A$ whose vector $\lambda(A)$ of eigenvalues (in, say, non-increasing order) belongs to $S$, and such sets are often called spectral convex sets. In this context, Kostant’s Convexity Theorem reduces to the Schur–Horn Theorem, which implies this special case of Proposition 4.3, see [SS20, Lemma 2.1].
@@ 5. Lifting convex algebro-geometric properties from a section
In view of Proposition 4.3, it is natural to ask which classes of convex subsets of Euclidean spaces are preserved under the bijection (1.1). In this section, we answer this question affirmatively for convex semi-algebraic sets (5.1), spectrahedral shadows (5.2), and rigidly convex sets (5.5), which proves Theorem B in the Introduction. Furthermore, we conjecture that spectrahedra are also preserved (5.4) and so are sums of squares (5.6).
@@
where $n=\dim\mathfrak{a}$, the $\rho^{j}\colon\mathfrak{g}_{\mathds{C}}\to\operatorname{End}(V_{j})$ are certain irreducible (complex) representations of the complex Lie algebra $\mathfrak{g}_{\mathds{C}}$, and $c_{j}(x)$ denotes the largest eigenvalue of $\rho^{j}(x)$. Note that for a certain choice of Hermitian inner product on $V_{j}$, the linear maps $\rho^{j}(z)$, for all $z\in V$, are self-adjoint (see [KS22, Lemma 4.6(a)]), so that the eigenvalues of $\rho^{j}(x)$ are all real, and hence “$c_{j}(x)\cdot\operatorname{Id}_{V_{j}}\succeq\rho^{j}(y)$” makes sense.
@@
5.2 has been proved in the special case of Example 4.4 in [SS20, Thm. 4.1].
@@
We note that 5.4 is still open even in the case of Example 4.4. One piece of evidence in this special case is [SS20, Thm. 3.3], which states that if $S$ is a symmetric polytope, then the spectral convex set $\mathsf{K}\cdot S$ is a spectrahedron; further progress was made in [Kum21]. Moreover, recall that all faces of a spectrahedron are exposed, and the bijection (1.1) preserves the class of convex sets all of whose faces are exposed [BGH13, BGH16].
Next, we show that the Generalized Lax Conjecture 2.15 implies 5.4:
@@
As in previous cases, the property of being rigidly convex is clearly preserved by intersecting with a vector subspace, so we shall only consider the map $S\mapsto\mathsf{K}\cdot S$. Let $S\subset\mathfrak{a}$ be a $W$-invariant rigidly convex set and $p$ be its defining polynomial as algebraic interior. Since $S$ is $W$-invariant and the defining polynomial is unique by Lemma 2.5, $p$ is also $W$-invariant. By the Chevalley Restriction Theorem, there exists a unique $\mathsf{K}$-invariant polynomial $\widetilde{p}\in{\mathds{R}}[V]^{\mathsf{K}}$ such that $\widetilde{p}|_{\mathfrak{a}}=p$. By 2.11 (i), we have that $p>0$ in the interior of $S$, and hence $\widetilde{p}>0$ in the interior of $\mathsf{K}\cdot S$. Moreover, $\widetilde{p}$ vanishes on $\partial(\mathsf{K}\cdot S)$ and has minimal degree among such polynomials because otherwise $p$ would not have minimal degree as the defining polynomial for $S$. Thus, $\mathsf{K}\cdot S$ is an algebraic interior with defining polynomial $\widetilde{p}$. It remains to show that $\widetilde{p}$ is real zero with respect to some (hence every) point in the interior of $\mathsf{K}\cdot S$. Without loss of generality, we may assume the origin $0\in\mathfrak{a}\subset V$ is in the interior of $S$, and hence $\widetilde{p}(0)=p(0)>0$. Given $v\in V$, let $g\in\mathsf{K}$ be such that $g\cdot v\in\mathfrak{a}$. Since $p$ is real zero with respect to $0\in\mathfrak{a}$, the univariate polynomial $t\mapsto p(t\,g\cdot v)$ has only real zeros. Since $\widetilde{p}(t\,v)=p(t\,g\cdot v)$, it follows that $t\mapsto\widetilde{p}(t\,v)$ has only real zeros. Thus, $\mathsf{K}\cdot S$ is rigidly convex. ∎
@@ 6. Optimization symmetry reduction with polar representations
One may hope the above reduction would enlarge the class of convex optimization problems that can be solved with SDP, by replacing (6.1) with (6.2). However, by 5.2, the convex set $\mathcal{O}\cap\mathfrak{a}$ is a spectrahedral shadow, i.e., (6.2) can be solved with SDP, if and only if $\mathcal{O}\subset V$ is a spectrahedral shadow, i.e., the original problem (6.1) can itself be solved with SDP. Nevertheless, the reduced problem (6.2) has $\operatorname{codim}\mathfrak{a}$ fewer variables, so, in practice, some efficiency is still gained.
@@
Let $\mathsf{K}=\mathsf{SO}(n)$, $n\geq 2$, act by conjugation on $V=\operatorname{Sym}^{2}({\mathds{R}}^{n})$, and recall this is a polar representation (4.4), with section $\mathfrak{a}\cong{\mathds{R}}^{n}$ given by the subspace of diagonal matrices. Let $\mathcal{O}\subset V$ be an $\mathsf{SO}(n)$-invariant convex set, fix $A\in\operatorname{Sym}^{2}({\mathds{R}}^{n})$ and choose $g\in\mathsf{SO}(n)$ such that $D=g^{-1}Ag$ is diagonal. Then,
@@
by 6.1, where the optimization problem on the left-hand side has $\frac{n(n+1)}{2}$ variables, and the one on the right-hand side has $n$ variables. Note that this polar representation is not irreducible, namely $\operatorname{Sym}^{2}({\mathds{R}}^{n})\cong{\mathds{R}}\langle\operatorname{I}_{n}\rangle\oplus\operatorname{Sym}^{2}_{0}({\mathds{R}}^{n})$, but this is irrelevant for the above symmetry reduction.
@@
by 6.1, where the optimization problem on the left-hand side has $\binom{n}{2}$ variables, and the one on the right-hand side has $\lfloor n/2\rfloor$ variables.
@@
6.1 applies mutatis mutandis to $\mathsf{K}$-representations $V$ of higher copolarity, since the orthogonal projection $\Pi$ onto a generalized section $\mathfrak{a}$ also satisfies $\Pi(\mathcal{O})=\mathcal{O}\cap\mathfrak{a}$, see [Men22].
@@
Even if a $\mathsf{K}$-representation $V$ has a vector subspace $\mathfrak{a}\subset V$ that meets all $\mathsf{K}$-orbits, 6.1 may fail if the assumption that $V$ is polar is dropped. Namely, consider the Hopf action of $\mathsf{U}(1)$ on $V={\mathds{C}}^{2}$ given by $e^{it}\cdot(z_{1},z_{2})=(e^{it}z_{1},e^{it}z_{2})$, and note that the vector subspace $\mathfrak{a}\cong{\mathds{R}}^{3}$ consisting of $(z_{1},z_{2})\in{\mathds{C}}^{2}$ with $\operatorname{Im}z_{2}=0$ meets all $\mathsf{U}(1)$-orbits (but not orthogonally). Although
@@
where $\Pi\colon{\mathds{C}}^{2}\to\mathfrak{a}$ is the orthogonal projection, in general $\Pi(\mathcal{O})\supsetneq\mathcal{O}\cap\mathfrak{a}$. For example, the convex hull $\mathcal{O}$ of the $\mathsf{U}(1)$-orbit of $(1,1)\in{\mathds{C}}^{2}$ is such that $\mathcal{O}\cap\mathfrak{a}$ is the straight line segment with endpoints ±$\pm(1,1)\in{\mathds{C}}^{2}$, while $\Pi(\mathcal{O})$ is the convex hull of the ellipse $\{(e^{it},\cos t)\in{\mathds{C}}^{2}:t\in{\mathds{R}}\}$. Note that the complexity of finding $\Pi(\mathcal{O})$ is the same as the original optimization problem, so nothing is gained by using (6.3).

### v2

@@ 3. Equivariant descriptions of invariant spectrahedra
Remark 3.10.
Starting from a spectrahedral description of size $d$ and $n$ variables, that is, with an affine-linear map $M\colon{\mathds{R}}^{n}\to\operatorname{Sym}^{2}({\mathds{R}}^{d})$, the equivariant spectrahedral description $\overline{M}\colon{\mathds{R}}^{n}\to\operatorname{Sym}^{2}({\mathds{R}}^{\overline{d}})$ obtained via Theorem A has size $\overline{d}\leq\binom{n+d-1}{n}$. This bound follows from the proof of Proposition 3.5, as $B(g)$ has $m=\dim W$ columns and $W\subset{\mathds{R}}[x_{1},\dots,x_{n}]_{<d}$. Although this bound is sharp in Example 3.7, the actual size $\overline{d}$ can be significantly smaller, e.g., in Example 3.8, where $\overline{d}=3$ and $\binom{n+d-1}{n}=21$, and in Example 3.9, where $\overline{d}=4$ and $\binom{n+d-1}{n}=15$.
As illustrated by Example 3.4, one cannot expect that $S_{\mathsf{G}}$ is a spectrahedron, even if $S$ is. The following partial converse to Proposition 3.5 shows that if $S_{\mathsf{G}}$ is a spectrahedron, which, moreover, is given by linear matrix inequalities coming from Construction 3.2, then, in fact, $S$ was already $\mathsf{G}$-invariant (assuming $\mathsf{G}$ is connected, cf. Remark 3.6).
@@
By Proposition 3.3, we have that $S_{\mathsf{G}}=\{x\in{\mathds{R}}^{n}:\widetilde{M}(x)\succeq 0\}$, so it suffices to show that $S_{\mathsf{G}}=S$, see (3.2). By Proposition 2.14 and Corollary 2.13, it suffices to prove $\partial S_{\mathsf{G}}\subset\partial S$. Note that $S_{\mathsf{G}}$ is a spectrahedron, since we are assuming $S_{\mathsf{G}}=\{x\in{\mathds{R}}^{n}:\overline{M}(x)\succeq 0\}$. Thus, given $x\in\partial S_{\mathsf{G}}$, the kernel of $\overline{M}(x)$ is nonzero, so there exists $\varphi\neq 0$ in $V$ such that $\widetilde{M}(x)(\varphi,\varphi)=0$, i.e.,
@@
The statement of Proposition 3.11 becomes false if we drop the assumption that $\mathsf{G}$ is connected. Indeed, if $\mathsf{G}$ is finite, then $\mathcal{V}$} itself is finite-dimensional.
@@ 4. Preliminaries from polar representations
A convex subset $S$ of $\mathfrak{a}\cong{\mathds{R}}^{n}$ that is invariant under $W=S_{n}$ is commonly called symmetric. The corresponding $\mathsf{O}(n)$-invariant convex subset of $V$ is given by the set of symmetric matrices $A$ whose vector $\lambda(A)$ of eigenvalues (in, say, non-increasing order) belongs to $S$, and such sets are often called spectral convex sets. In this context, Kostant’s Convexity Theorem reduces to the Schur–Horn Theorem, which implies this special case of Proposition 4.3, see [SS25, Lemma 2.1].
@@ 5. Lifting convex algebro-geometric properties from a section
In view of Proposition 4.3, it is natural to ask which classes of convex subsets of Euclidean spaces are preserved under the bijection (1.1). In this section, we answer this question affirmatively for convex semi-algebraic sets (Proposition 5.1), spectrahedral shadows (Proposition 5.2), and rigidly convex sets (Proposition 5.5), which proves Theorem B in the Introduction. Furthermore, we conjecture that spectrahedra are also preserved (Conjecture 5.4) and so are sums of squares (Conjecture 5.6).
@@
where $n=\dim\mathfrak{a}$, the $\rho^{j}\colon\mathfrak{g}_{\mathds{C}}\to\operatorname{End}(V_{j})$ are certain irreducible (complex) representations of the complex Lie algebra $\mathfrak{g}_{\mathds{C}}$, and $c_{j}(x)$ denotes the largest eigenvalue of $\rho^{j}(x)$. Note that for a certain choice of Hermitian inner product on $V_{j}$, the linear maps $\rho^{j}(z)$, for all $z\in V$, are self-adjoint (see [KS22, Lemma 4.6(a)]), so that the eigenvalues of $\rho^{j}(x)$ are all real, and hence $c_{j}(x)\cdot\operatorname{Id}_{V_{j}}\succeq\rho^{j}(y)$” makes sense.
@@
Proposition 5.2 has been proved in the special case of Example 4.4 in [SS25, Thm. 4.1].
@@
We note that Conjecture 5.4 is still open even in the case of Example 4.4. One piece of evidence in this special case is [SS25, Thm. 3.3], which states that if $S$ is a symmetric polytope, then the spectral convex set $\mathsf{K}\cdot S$ is a spectrahedron; further progress was made in [Kum21]. Moreover, recall that all faces of a spectrahedron are exposed, and the bijection (1.1) preserves the class of convex sets all of whose faces are exposed [BGH13, BGH16].
Next, we show that the Generalized Lax Conjecture 2.15 implies Conjecture 5.4:
@@
As in previous cases, the property of being rigidly convex is clearly preserved by intersecting with a vector subspace, so we shall only consider the map $S\mapsto\mathsf{K}\cdot S$. Let $S\subset\mathfrak{a}$ be a $W$-invariant rigidly convex set and $p$ be its defining polynomial as algebraic interior. Since $S$ is $W$-invariant and the defining polynomial is unique by Lemma 2.5, $p$ is also $W$-invariant. By the Chevalley Restriction Theorem, there exists a unique $\mathsf{K}$-invariant polynomial $\widetilde{p}\in{\mathds{R}}[V]^{\mathsf{K}}$ such that $\widetilde{p}|_{\mathfrak{a}}=p$. By Proposition 2.11 (i), we have that $p>0$ in the interior of $S$, and hence $\widetilde{p}>0$ in the interior of $\mathsf{K}\cdot S$. Moreover, $\widetilde{p}$ vanishes on $\partial(\mathsf{K}\cdot S)$ and has minimal degree among such polynomials because otherwise $p$ would not have minimal degree as the defining polynomial for $S$. Thus, $\mathsf{K}\cdot S$ is an algebraic interior with defining polynomial $\widetilde{p}$. It remains to show that $\widetilde{p}$ is real zero with respect to some (hence every) point in the interior of $\mathsf{K}\cdot S$. Without loss of generality, we may assume the origin $0\in\mathfrak{a}\subset V$ is in the interior of $S$, and hence $\widetilde{p}(0)=p(0)>0$. Given $v\in V$, let $g\in\mathsf{K}$ be such that $g\cdot v\in\mathfrak{a}$. Since $p$ is real zero with respect to $0\in\mathfrak{a}$, the univariate polynomial $t\mapsto p(t\,g\cdot v)$ has only real zeros. Since $\widetilde{p}(t\,v)=p(t\,g\cdot v)$, it follows that $t\mapsto\widetilde{p}(t\,v)$ has only real zeros. Thus, $\mathsf{K}\cdot S$ is rigidly convex. ∎
@@ 6. Optimization symmetry reduction with polar representations
Let $\mathsf{K}=\mathsf{SO}(n)$, $n\geq 2$, act by conjugation on $V=\operatorname{Sym}^{2}({\mathds{R}}^{n})$, and recall this is a polar representation (Example 4.4), with section $\mathfrak{a}\cong{\mathds{R}}^{n}$ given by the subspace of diagonal matrices. Let $\mathcal{O}\subset V$ be an $\mathsf{SO}(n)$-invariant convex set, fix $A\in\operatorname{Sym}^{2}({\mathds{R}}^{n})$ and choose $g\in\mathsf{SO}(n)$ such that $D=g^{-1}Ag$ is diagonal. Then,
@@
by Symmetry Reduction 6.1, where the optimization problem on the left-hand side has $\frac{n(n+1)}{2}$ variables, and the one on the right-hand side has $n$ variables. Note that this polar representation is not irreducible, namely $\operatorname{Sym}^{2}({\mathds{R}}^{n})\cong{\mathds{R}}\langle\operatorname{I}_{n}\rangle\oplus\operatorname{Sym}^{2}_{0}({\mathds{R}}^{n})$, but this is irrelevant for the above symmetry reduction.
@@
by Symmetry Reduction 6.1, where the optimization problem on the left-hand side has $\binom{n}{2}$ variables, and the one on the right-hand side has $\lfloor n/2\rfloor$ variables.
@@
One may hope that Symmetry Reduction 6.1 would enlarge the class of convex optimization problems that can be solved with SDP, by replacing (6.1) with (6.2). However, by Proposition 5.2, the convex set $\mathcal{O}\cap\mathfrak{a}$ is a spectrahedral shadow, i.e., (6.2) can be solved with SDP, if and only if $\mathcal{O}\subset V$ is a spectrahedral shadow, i.e., the original problem (6.1) can itself be solved with SDP. Nevertheless, some efficiency is gained, as the number of variables in the reduced problem (6.2) is $\dim\mathfrak{a}$, while the original problem has $\dim V$ variables. As discussed in Section 4, polar representations $V$ of connected groups $\mathsf{K}$ are orbit-equivalent to the isotropy representation of a symmetric space. In this language, $\dim V$ is the dimension of the symmetric space, and $\dim\mathfrak{a}$ is its rank. If $V$ is an irreducible $\mathsf{K}$-representation, it follows from the classification of symmetric spaces (see [Hel78, Chap. X]) that $\dim\mathfrak{a}=O(\sqrt{\dim V})$. However, $\dim\mathfrak{a}$ may be larger if irreducibility is dropped; e.g., the natural action of $\mathsf{K}=\mathsf{SO}(2)^{k}$ on $V={\mathds{R}}^{2k}$ is polar with section $\mathfrak{a}={\mathds{R}}^{k}$.
@@
Symmetry Reduction 6.1 applies mutatis mutandis to $\mathsf{K}$-representations $V$ of higher copolarity, since the orthogonal projection $\Pi$ onto a generalized section $\mathfrak{a}$ also satisfies $\Pi(\mathcal{O})=\mathcal{O}\cap\mathfrak{a}$, see [Men22].
@@
Even if a $\mathsf{K}$-representation $V$ has a vector subspace $\mathfrak{a}\subset V$ that meets all $\mathsf{K}$-orbits, Symmetry Reduction 6.1 may fail if the assumption that $V$ is polar is dropped. Namely, consider the Hopf action of $\mathsf{U}(1)$ on $V={\mathds{C}}^{2}$ given by $e^{it}\cdot(z_{1},z_{2})=(e^{it}z_{1},e^{it}z_{2})$, and note that the vector subspace $\mathfrak{a}\cong{\mathds{R}}^{3}$ consisting of $(z_{1},z_{2})\in{\mathds{C}}^{2}$ with $\operatorname{Im}z_{2}=0$ meets all $\mathsf{U}(1)$-orbits (but not orthogonally). Although
@@
where $\Pi\colon{\mathds{C}}^{2}\to\mathfrak{a}$ is the orthogonal projection, in general $\Pi(\mathcal{O})\supsetneq\mathcal{O}\cap\mathfrak{a}$. For example, the convex hull $\mathcal{O}$ of the $\mathsf{U}(1)$-orbit of $(1,1)\in{\mathds{C}}^{2}$ is such that $\mathcal{O}\cap\mathfrak{a}$ is the straight line segment with endpoints $\pm(1,1)\in{\mathds{C}}^{2}$, while $\Pi(\mathcal{O})$ is the convex hull of the ellipse $\{(e^{it},\cos t)\in{\mathds{C}}^{2}:t\in{\mathds{R}}\}$. Note that the complexity of finding $\Pi(\mathcal{O})$ is the same as the original optimization problem, so nothing is gained by using (6.3).
Remark 6.5.
To implement Symmetry Reduction 6.1, one must find $g\in\mathsf{K}$ such that $g^{-1}\cdot v$ belongs to $\mathfrak{a}\subset V$, or, alternatively, find a section $\mathfrak{a}^{\prime}\subset V$ with $v\in\mathfrak{a}^{\prime}$. For generic $v\in V$, the latter can be achieved by setting $\mathfrak{a}^{\prime}$ to be the orthogonal complement of the tangent space at $v$ to the $\mathsf{K}$-orbit through $v$. In practice, this amounts to performing matrix multiplications and solving a linear systems of equations. However, in some situations (such as Example 6.2), the former problem reduces to diagonalizing a symmetric matrix, for which faster methods are available.